Volume 1, Number 4 December 1992

TEX and TUG NEWS

for and by the TEX community

A Publication of the TEX Users Group Electronic version TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives, as well as at the heidelberg, labrea, and ymir archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font selection scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@math.ams.org with the subject line NEWSLETTER. Regular mail should be addressed to the Publications Committee, c/o TUG Office at the address below.

TEX Users Group

P.O. Box 869  Phone: 805-899-4673
Santa Barbara, CA 93102 USA E-mail: tug@math.ams.org

TEX is a trademark of the American Mathematical Society.

TEX and TUG NEWS (ISSN 1065-240X) is published four times a year, February, July, October and December, by the TEX Users Group, 735 State Street, P.O. Box 869, Santa Barbara, CA 93102, U.S.A. The 1992 dues for individual members are $60.00 for regular members and $50.00 for students.

Second-class postage paid at Providence, RI and additional mailing offices. Postmaster: Send address changes to the TEX Users Group, P.O. Box 869, Santa Barbara, CA 93102, U.S.A.

Membership in the TEX Users Group is for the calendar year, and includes all issues of TUGboat and TEX and TUG NEWS for the year in which membership begins or is renewed. Individual membership is open only to named individuals, and carries with it such rights and responsibilities as voting in the annual election.

Editorial

Well, this is it. Our first full year as a newsletter is complete. Hopefully, you’ve learned a few new tricks, found out about happenings around the TEX world, and generally enjoyed what we’ve published. I must say, though, that reader response has been rather disappointing, in that few comments have come back. Some would say that means we aren’t offending anyone’s sensibilities — others would say it’s a sign of quiet indifference! ’nuff said. Door’s open …

Feature article this month is “TEX Conventions Concerning Languages”, by Yannis Haralambous (pp. 3–10). It’s the first time we’ve had to compromise on portability: the font table on pp. 4–5 requires the use of dcr10. But it seemed an acceptable break this once, and of course provided us the opportunity of then turning a whole slew of questions on Yannis about just what DC fonts are! See the companion piece: “DC fonts: questions and answers” (part 1: pp. 15–17). If we didn’t ask the questions you need answers for, send them in and we’ll select the most representative ones for another Q and A.

Lots of news, as usual, and something which warrants some serious thought: electronic documents and the law. GUTenberg is organising a one-day meeting on the subject next spring, and is collecting topics now; see p. 19. All the regular columns are here … and we may add a few new ones next year. Stay tuned.

To round out the year, we’ve put together an Index of all issues since the prototype, including the contents of this issue. Go use it and see if it helps locate your favourite pieces. Then write to me and tell me what’s missing!

And some major developments within TUG: a new Executive Director has been named. A TUG member since 1984, Pat Monohon is also an experienced TEX user. Pat is just now in the process of settling in — and operations are moving to Santa Barbara. While our traditional home has been in Providence since TUG’s inception, the move is expected to be smooth and not impede or interrupt things unduly. Welcome aboard, Pat! See p. 22 for details.

With Pat’s arrival, we also see the departure of Ron Whitney. I’ve been working with Ron for the past few years not only as a board member, but also in connection with organising some of TUG’s annual meetings and producing the attendant proceedings. He has shown a great deal of commitment and support for TUG even when our recent bumpy past has made his job very difficult at times. I would like to recognise his efforts of the past few years. Good luck in your new adventures, Ron. And good fortune.

The other news is that, in the absence of nominations for president, the Executive Committee had to find a pragmatic alternative solution, which is outlined in a piece by Malcolm Clark, TUG’s outgoing president (p. 22).

Last issue I suggested that people who provide support for users might find it useful to exchange ideas. Another area for exchange concerns interesting symbols (for mathematics, physics, linguistics [had to throw that one in! ] ) which people have to cobble together — the Cauchy Integral was one we had. If someone wants to take this on as an occasional column, please get in touch with me.

And I might as well toss out another idea. We’re in a transition period, it seems to me, between the basic TEX building blocks and a growing number of enhancements, additions, and improvements. Some users have all the latest stuff, some have only parts, some have changed nothing. There’s TEX v.3.14 now out there but not on all machines; there’s the New Font Selection Scheme (NFSS); there are DC fonts and virtual fonts; and coming soon — L^{A}TEX3, just to mention a few. Almost old news now (but certainly not standardised) is the integration of PostScript into TEX: graphics, non-CM fonts, images, and so on.

But … what if all you know is “the basic TEX building blocks”? What if you need to get the NFSS set up on your PC or Mac or UNIX box or whatever? Should you worry about virtual fonts? How do you get at those PostScript fonts? Or even, how do you get at the PostScript versions of the standard CM fonts? All these questions! Where are the answers? The solutions? The warnings?!

Ask yourself those questions. Do you know the answers? Do you know where to find the answers? Are you worried?! How can TTN play a role here? What part could you play in this? Do you have a particular expertise which could be parlayed into providing advice? And what are your questions?

While many people have access to colleagues, newsgroups and electronic mail, many more TEX users and TUG members don’t have those avenues of information. They rely on publications, on read.me files, on friends over in the engineering department (you never know!). Can we take these issues, and bring some order and sense to them on the pages of TTN? Or perhaps through a combined effort with both TTN and TUGboat?

That’s our challenge: not to simply reflect our users, but to enjoin them to participate, to share their expertise. Remember to renew for next year, otherwise you won’t learn the answers to these and other burning questions in the world of TEX and TEX users!

Christina Thiele

Editor, TEX and TUG NEWS

Did you know …?

… the student rate has now been set at 50% of the regular individual membership rate? That means students pay $30.00 US for 4 issues of TUGboat, 4 issues of TEX and TUG NEWS, and a membership directory, as well as discounts on items purchased from TUG, and on TUG courses and the annual meeting. A bargain, for sure!

TEX Conventions Concerning Languages

Yannis Haralambous

Chair, Technical Working Group^{1} on Multiple Language Coordination

(WG-92-03)^{2}

… each language draws a magic circle round the people to which it belongs, a circle from which there is no escape save by stepping out of it into another. Marshall McLuhan,^{3} The Gutenberg Galaxy (1962)

What are DC Fonts?

When I first started using TEX to typeset my thesis (in French), I noticed that it was producing far too many black overfull boxes. Trying to find out what was going wrong, I realized that TEX (actually good old Textures v1.0) wasn’t very cooperative in hyphenating; as a matter of fact, some words were not hyphenated at all. Some further experimentation revealed the fact that words containing accents were never hyphenated! A deeper search in the TEXbook clarified the situation: the \accent primitive, included in the expansion of accent macros, was introducing an explicit kern, which in turn aborted the hyphenation process for the particular word.

So it was necessary to use accented characters, which are quite common in European languages, as a whole, rather than merely accenting them by use of the \accent primitive. This has finally become possible with the release of TEX 3, which allows 8-bit input, solving the problem in situations where one wants to use fonts other than Computer Modern. Still, a standard 8-bit “Computer Modernlike” font was needed, to ensure TEX’s tradition of portability and structural consistency. Jan Michael Rynning and Norbert Schwarz undertook the task of establishing a new font table. Since the Cork Meeting, in 1990, TUG and the TEX community have adopted this font table as a new standard for TEX, known as “DC encoding”, “Cork encoding”, or “extended TEX font encoding” (see TUGboat 11, no. 4, pp. 514–516). Norbert Schwarz has written the - code for a family of Computer Modern-like fonts with this table, called DC fonts.

1I would like to thank Johannes Braams, Mike Ferguson, Bernard Gaulle, Jörg Knappen, Jan Michael Rynning, Rainer Schöpf and Erik-Jan Vens for their contributions.

2See pp. 7–8 in previous issue of TTN for description of this Working Group –Ed.

3Actually quoting Wilhelm von Humboldt, and his “strategy any culture must resort to in a period like this … ”

It might be appropriate to mention that Norbert does not consider the - code of these fonts to be absolutely finished — small corrections on the shapes of characters may still occur; nevertheless, the encoding scheme will not change, and the TEX community is invited to begin migrating from the Computer Modern encoding described in the five volumes of Computers and Typesetting to the new DC Font Encoding, shown below:

          Font dcr10

’0 ’1 ’2 ’3 ’4 ’5 ’6 ’7
’00x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ
’01x ˘  ¯  ˙  ¸  ˛  ‚ ‹ ›
’02x “ ” „ « » – —
’03x � ı ȷ ff fi fl ffi ffl
’04x ␣ ! " # $ % & ’
’05x ( ) * + , - . /
’06x 0 1 2 3 4 5 6 7
’07x 8 9 : ; < = > ?
’10x @ A B C D E F G
’11x H I J K L M N O
’12x P Q R S T U V W
’13x X Y Z [ \ ] ^_
’14x ‘ a b c d e f g
’15x h i j k l m n o
’16x p q r s t u v w
’17x x y z { | } ~ -
’20x Ă Ą Ć Č Ď Ě Ę Ğ
’21x Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
’22x Ř Ś Š Ş Ť Ţ Ű Ů
’23x Ÿ Ź Ž Ż Ĳ İ đ §
’24x ă ą ć č ď ě ę ğ
’25x ĺ ľ ł ń ň ŋ ő ŕ
’26x ř ś š ş ť ţ ű ů
’27x ÿ ź ž ż ĳ ¡ ¿ £

’30x À Á Â Ã Ä Å Æ Ç
’31x È É Ê Ë Ì Í Î Ï
’32x Ð Ñ Ò Ó Ô Õ Ö Œ
’33x Ø Ù Ú Û Ü Ý Þ SS
’34x à á â ã ä å æ ç
’35x è é ê ë ì í î ï
’36x ð ñ ò ó ô õ ö œ
’37x ø ù ú û ü ý þ ß

This migration will consist of:

1. adding the new fonts to the system (the Computer Modern fonts are still necessary because the Greek uppercase letters are not part of the DC text font encoding — they will be included in the forthcoming DC math fonts);

2. replacing plain.tex, lplain.tex by appropriate new versions (Rainer Schöpf reports that the NFSS is already fully compatible with DC fonts), so you just have to be sure to get the latest version (v.1.0a, with last bug fixed Sept. 20, 1992) — and the DC fonts will work;

3. taking care, when writing TEX code, to use high-level macros for certain purposes involving character encoding: \`a, \oe, $\Sigma$ will always produce the correct results: à, œ, Σ. The apparently equivalent code

\accent'022a, \char'033, \char"06 will produce à, œ, Σ in a CM font

environment but „a, ff, ˚ in a DC font environment.

Recommendations on certain macro and file names concerning languages

A few words

All macro commands used in the plain.tex format are defined and explained in the TEXbook by Knuth, the very author of TEX. These macros are a standard part of TEX and this fact guarantees the portability of TEX. The DC fonts and TEX 3.xx’s language switching features require new macros, which will also have to be standardized.

These macros are of two kinds:

1. macros for accessing accented or special characters which are not available in the Computer Modern fonts (see “Macros for accented and special characters”, below), and cannot be obtained by the standard plain.tex accenting macros (of course, any character can be accessed by using, for example, the \char macro, but this does not really optimize readability of TEX code);

2. macros for language switching (see “Language-switching macros”, below). Since it is now possible to typeset a multilingual text where each language uses its own hyphenation rules, its own fonts and eventually its own direction of script, there must be a standard (and simple) way to switch between these languages.

So if, for example, you want to include German or French words in your English text — “Wagner’s Götterdämmerung was very appreciated by the 19th century’s bourgeoisie” — they will have to be hyphenated according to German and French rules. To indicate this to TEX, macros have to be selected which are easy to remember, short, and universally acceptable.

For this, the standard has basically been taken from the standard 2-letter ISO 639 language codes^{4} as control sequence names (see “Languageswitching macros”, below).

The previous example is then written:

   \US Wagner's \DE {\it G\"otterd\"ammerung}
   \US was very appreciated by the 19th century's
   \FR {\it bourgeoisie}

3. Finally, besides macro names, the files containing hyphenation patterns should have standardized names and the most natural choice would be to take the 2-letter macro name defined in the section “Language-switching macros” as a (lowercase) prefix to the usual “hyphen.tex”. In this way, Frank Liang’s file of American English hyphenation patterns would become ushyphen.tex, while there is already a vast collection of other hyphenation pattern files available.^{5}

Macros for accented and special characters

Besides the accent and special character macros defined in the TEXbook (pp. 52,

339) the following macros are recommended:

4With the following exceptions: TY for Thai (instead of TH) and FF for Afar (instead of AA); US and GB are chosen for American and British English, respectively; SB is chosen for Sorbian and SE for Sami. The codes LA (Latin), SA (Sanskrit), EN (English), and SH (Serbocroatian)

are not used.

5At the present time, hyphenation pattern files exist for Armenian, Bulgarian (β), Catalan (β), Croatian, Czech, Danish, Dutch, Esperanto, Estonian, Finnish, French, German, Greek (modern and ancient), Hungarian, Icelandic, Italian, Kirundi (β), Latin, Lithuanian, Norwegian, Polish, Portuguese (β), Russian, Slovak, Swedish, UK English, and US English; work is in progress on Cambodian, Swahili, Yiddish, and Yoruba.

Type  to get (Description)
\r u  ů  (Czech circle accent)
\k e  ę  (Polish ogonek subscript)
\v d \v D ď Ď (Czech d D with haček)
\v t \v T ť Ť (Czech t T with haček)
\v l \v L ľ Ľ (Slovakian l L with haček)
\th \TH  þ Þ (Icelandic thorn)
\dh \DH  ð Ð (Icelandic eth)
\dj \DJ  đ Ð (Serbocroatian dj)
\ng \NG  ŋ Ŋ (Sami ng)
<< >>  « » (French guillemets)
,, ``  ,, “ (German Gänsefüßchen)

Notes

• The Dutch character ĳ Ĳ, as in:

Blĳf, wĳl ’k mĳn tĳd Blĳ ’t ĳ-rĳm wĳd, Blĳf, blĳf mĳ bĳ, gĳ Ĳ, Stĳf vrĳ, wĳl ’k lĳm, Mĳn ĳ-rĳk rĳm; Mĳn wĳs, mĳn prins zĳt gĳ

is a ligature (just like ffi, ffl) and should be accessed as such; special Dutch (virtual) fonts could eventually provide this ligature and the \NL macro would in that case automatically switch to these fonts.

• The Scandinavian character å Å can be obtained either by the standard

plain.tex macros \aa \AA or by \r a \r A.

• The macro \c serves as well for the French, Catalonian and Latvian cedilla as for the Latvian letter ‘g, with the diacritic above the letter: \c g for ‘‘g’.

• The Turkish letters ı I, i İ are obtained by the usual \i I, i \.I.

• The French and German quotation marks « » ,, “ do not require macros! They are obtained as ligatures (just like the English quotation marks “ ”).^{6}

6Ligatures are used to simply typeset the French and German quotation marks as any other character. If the use of quotation marks affects the global aspect of the page, the use of macros is recommended (as in french.sty by Bernard Gaulle, in the case of French traditional typography: quotations require a left guillemet at the beginning of every paragraph and secondorder quotations require one at the beginning of every line).

8 TEX and TUG NEWS Vol. 1, No. 4, 1992

Language-switching macros

The following table indicates two-letter control sequences and the corresponding languages:

Two-Letter Macros for Language Switching

Abkhazian  \AB Frisian  \FY Malagasy  \MG Sorbian  \SB
Afan Oromo \OM Galician  \GL Malay  \MS Spanish  \ES
Afar  \FF Georgian  \KA Malayalam  \ML Sudanese  \SU
Afrikaans  \AF German  \DE Maltese  \MT Swahili  \SW
Albanian  \SQ Greek  \EL Maori  \MI Swedish  \SV
Amharic  \AM Greenlandic  \KL Marathi  \MR Tagalog  \TL
Arabic  \AR Guarani  \GN Moldavian  \MO Tajik  \TG
Armenian  \HY Gujarati  \GU Nauru  \NA Tamil  \TA
Assamese  \AS Hausa  \HA Nepali  \NE Tatar  \TT
Aymara  \AY Hebrew  \HE Norwegian  \NO Telugu  \TE
Azerbaijani  \AZ Hindi  \HI Occitan  \OC Thai  \TY
Bashkir  \BA Hungarian  \HU Oriya  \OR Tibetan  \BO
Basque  \EU Icelandic  \IS Pashto  \PS Tigrinya  \TI
Bengali  \BN Indonesian  \ID Persian  \FA Tonda  \TO
Bhutani  \DZ Interlingua  \IA Polish  \PL Tsonga  \TS
Bihari  \BH Interlingue  \IE Portuguese  \PT Turkish  \TR
Bislama  \BI Inuktitut  \IU Punjabi  \PA Turkmen  \TK
Breton  \BR Inupiak  \IK Quechua  \QU Twi  \TW
Bulgarian  \BG Irish  \GA Rhaeto-Rom. \RM Uigur  \UG
Burmese  \MY Italian  \IT Romanian  \RO UK English \GB
Byelorussian \BE Japanese  \JA Russian  \RU Ukrainian  \UK
Cambodian \KM Javanese  \JW Sami  \SE Urdu  \UR
Catalan  \CA Kannada  \KN Samoan  \SM US English \US
Chinese  \ZH Kashmiri  \KS Sangho  \SG Uzbek  \UZ
Corsican  \CO Kazakh  \KK Scots Gaelic \GD Vietnamese \VI
Croatian  \HR Kinyarwanda \RW Serbian  \SR Volapuk  \VO
Czech  \CS Kirghiz  \KY Sesotho  \ST Welsh  \CY
Danish  \DA Kirundi  \RN Setswana  \TN Wolof  \WO
Dutch  \NL Korean  \KO Shona  \SN Xhosa  \XH
Esperanto  \EO Kurdish  \KU Sindhi  \SD Yiddish  \YI
Estonian  \ET Laothian  \LO Singhalese  \SI Yoruba  \YO
Faroese  \FO Latvian  \LV Siswati  \SS Zhuang  \ZA
Fiji  \FJ Lingala  \LN Slovak  \SK Zulu  \ZU
Finnish  \FI Lithuanian  \LT Slovenian  \SL
French  \FR Macedonian  \MK Somali  \SO

These macros can have different expansions according to their context. Nevertheless, they should stay on a low level (in particular, they should be independent of any eventual macro package like Plain, L^{A}TEX, etc.) and there is a minimum number of features they strictly must provide:

1. a switch to the appropriate \language value and hence to the appropriate set of hyphenation patterns;

 2. the correct values of \lefthyphenmin and \righthyphenmin;

3. if the language requires a different set of fonts (either virtual or real), a switch to these fonts must be provided. In such a case, the macros for languages covered by the DC font encoding should switch back to the original DC-encoded fonts. The attribute codepage in the forthcoming NFSS v.2 illustrates this switch in the best possible way: the font’s family, shape, series and size remain the same, only the set of characters is changed;

4. if the language is written in a right-to-left direction, the appropriate TeX--XeT R-mode primitives should be included (in this case the macros for left-to-right languages should include the corresponding L-mode primitives).^{7}

The role of these macros is comparable to the one of macros such as \rm, \bf, etc. Although they can have different expansions (\bf has different expansions in Plain TEX and in NFSS), there is a consensus on their effect on the text; they can be — and frequently are — included in the expansions of other macros; and finally they are part of the first macros a TEX novice will learn and use.

Notes

• Languages have been classified according to these four criteria: (a) hyphenation patterns, (b) left- and righthyphenmin values, (c) the need for virtual fonts or use of a different script, and hence the need for different fonts, and (d) direction of script. So, for example, British English requires a set of hyphenation patterns different from those for American English; Lithuanian requires the use of virtual fonts; Greek and Russian use scripts other than Latin; Arabic and Hebrew are written from right-to-left; etc.

If further differentiation (based on these criteria) is needed, new macros will be added (for example, to eventually distinguish between Dutch and Flemish, German and Austrian or Schwitzer Dütsch, Brazilian and Portuguese, etc.) On the other hand, languages which present identical characteristics will use the same macro.

• Serbian and Croatian are actually a single language (Serbocroatian) with two main variants: Ekavian and Jekavian. Ekavian is mostly used in Serbia, while Jekavian is mostly used in Croatia; the former is mostly written in Cyrillic script, while the latter is mostly written in Latin script. In this spirit, the macros \SR and \HR should be understood as “Ekavian, written in Cyrillic script” and “Jekavian, written in Latin script”, respectively.

7TeX--XeT is Peter Breitenlohner’s implementation of TeX-XeT, the bi-directional version of TEX, described in TUGboat 8, no. 1, pp. 14–25. TeX--XeT produces a regular dvi file and hence can completely — and without any loss of compatibility — replace TEX, a practice that is strongly recommended.

• \HE should be understood as “Ivrit” or “Modern Hebrew”. It has the same script as Yiddish (\YI), but different hyphenation rules.

• Armenian transliteration and hyphenation patterns follow the Western pronunciation of this language (used by scholars and by the Armenian communities in Europe and the States). The use of virtual fonts is recommended for the input of Armenian transliterated according to the Eastern pronunciation.

• This list covers only languages and/or scripts used today. Older languages, such as Latin, Ancient Greek, Classical Hebrew, Aramaic, Syriac, Sanskrit, Ottoman Turkish, etc. are not covered, and the use of two-letter control sequences for them is not recommended. Please contact the Technical Working Group on Multiple Languages Coordination for further information [see TTN 1,3:7 –Ed].

Names of hyphenation pattern files

As previously mentioned, these should be named as

                      **hyphen.tex

where ** is the 2-letter language code outlined in “Language-switching macros”, above, in lowercase form. The standard American English hyphenation patterns file hyphen.tex will be called ushyphen.tex. Files for French, German, Dutch

will be called frhyphen.tex, dehyphen.tex, nlhyphen.tex, etc.

Correction notice: margin marks from TTN 1,3

Christina Thiele

In the last issue of TTN, Peter Schmitt’s “Hey, it works!” piece on margin marks was incomplete — the printers thought the examples of margin marks in the form of vertical rules were errors marring the left edge of the master copy. So, they very kindly and carefully removed them … Well, here they are, in their correct glory. Please accept my apologies for the error, and make a cross-reference note between p. 13 in TTN 1,3, and p. 10 in this issue. The correct output:

• The mark can be changed by \let\margmark=\bar even in the middle of a marked paragraph, provided one accepts that the line where the change takes places shows both marks (as shown to the left). If applicable, the definition of \margmark should use explicit fonts. Otherwise a font change in the text may influence the appearance of the mark.

New Publications^{1}

Rames Abdelhamid. Das Vieweg L^{A}TEX-Buch: Eine praxisorientierte Einführung. Braunschweig: Vieweg-Verlag, 1992. About 200pp. ISBN 3-528-05145-0.

• This book provides a well-structured and readable introduction to L^{A}TEX in combination with the german.sty option.

George Gratzer. Math into TEX. New York: Springer Verlag. 187pp. (softcover). Approx. $34.00 US. ISBN 0-8176-3637-4. [Avail. Dec. 1992]

• “This book is for the mathematician, engineer, or scientist who wants to write and typeset articles with mathematical formulas but who does not want to spend a great deal of time learning how to do it. It assumes little familiarity with TEX or L^{A}TEX.” [From promotional material]

Piotr Wyrostek, trans. L^{A}TEX: System przygotowywania dokumentöw. Przewodnik ürytkownika i prodrëcznik. Polish translation of Lamport’s L^{A}TEX manual. Krakow: Ariel. Price: 91000 zl. ISBN 83-900460-1-6.

Michaela Lichá and Oldřich Ulrych. AMS-TEX verze 2.1. Prague: Czechoslovak TEX Users Group, 1992. 80pp. [In Czech]

• “This publication provides a short (but complete) tour through AMS-TEX 2.1 together with amsppt.sty 2.1. Almost all features are demonstrated with a lot of examples. The book is organized in such a way that it is useful for novices (basic textbook), and for advanced users (quick reference quide) as well. Contains a table of all symbols.” [Info. from authors]

Latest issues of other TEX newsletters

Baskerville: vol. 2, no. 1, March 1992, 12pp. Official newsletter of the UK TEX

Users’ Group. Die TEXnische Komödie: vol. 4, no. 2, August 1992, 64pp. Official newsletter of

DANTE, Deutschsprachige Anwendervereinigung TEX e.V. MAPS: issue 92.2, 1992, 176pp. Official publication of NTG, Nederlandstalige

TEX Gebruikersgroep (MAPS = “Minutes and APendiceS”). TEXbulletin: no. 2, 1992, 48pp; no. 3, 1992, 64pp. Official newsletter of CsTUG.^{2} TUGboat: 13, no. 3, 1992, 162pp. Proceedings of TUG’92 meeting in Portland,

Oregon.

1NOTE: In the last issue of TTN, Peter Schmitt’s address suffered a typo. It should read a8131dal@awiuni11.EDVZ.UniVie.ac.at. I apologise for the inconvenience this has caused –Ed.

2Re: TTN 1,3 notice on the Prague Proceedings: Jiří Zlatuska of CsTUG asks that orders for the publication include a shipping and handling charge: $5 for surface mail, $10 for airmail in addition to the 30,– DM cost. –Ed.

Typographer’s Inn

Peter Flynn

University College Cork

                                          cbts8001@iruccvax.ucc.ie

Dropped caps

TTN’s editor, Christina Thiele, asked me about a problem which comes up at my work with regularity: dropped initial caps. I have a macro of my own to do this, and doubtless others have as well. I set the cap letter in whatever size is designed, and get TEX to measure its width, then set it, back up the following paragraph with the right negative \vskip, and then use \parshape to set the remainder of the para with two or three shortened lines at the top. This is a bit clumsy: has anyone a more elegant approach? I’ve looked at two or three macros which have come my way, and there seem to be several approaches, based on several different design decisions:

• How many lines do we want to dropindent? Most implementations assume you want to dropindent as many lines as the height of the cap dictates, but this may not always be possible: given a specific x-height and leading, you may have to choose your large cap size with care.

• Do we want the large cap to be aligned with the baseline of the lowest of the hanging indented lines it is inserted into? This looks the most elegant and would seem to be the obvious way to go.

• Do we instead want to insert the large cap into the first line only and have it stick up above the line, possibly quite a long way? This to my mind looks inelegant unless handled very carefully.

• Do we want the remainder of the first word in small caps, which was common practice until this century. It avoids the raggedness of having lowercase letters with uneven heights (because of the ascenders) following immediately after the large cap. You could also use ordinary caps, but they tend to obtrude in a line where the remaining words are lowercase.

As far as I know there is no way to get TEX or L^{A}TEX to wander the starts of the indented lines around the shape of the dropped initial. In visual dtp systems, this is not so much of a problem because the big letter is just a bunch of dots on the screen, or an outline which you can wander round. In TEX, a character is a box, and you can only trespass within it by manual intervention, which I’ve always thought a pity. Is there any way to do this in TEX if you are working with PostScript outlines, perhaps a preprocessor which, given the \baselineskip, would calculate how far leftwards to extend the start of each line so that it was a given distance away from the edge of the letter.

The smiley

I suppose I’ll have to give the source code for the smiley face I used last time. I don’t normally go in for these devices, but I do a lot of e-mail, and the use of the “smiley” (the sideways smiling face (:-) and its many variants) has become habitual. Typing the last column it dawned on me that that handy symbol was a short sort in CM, so I used the following kludge:

\font\tmi=cmmi7
\font\tsy=cmsy5
\font\bsy=cmsy10 scaled\magstep4
\newbox\sbox
\setbox\sbox=\vbox{\hsize=1pc\baselineskip=5pt
                   \centerline{{\tsy\char'014\ \char'014}}
                   \centerline{{\tsy\char'064}}\vskip-1pt
                   \centerline{{\tmi\char'136}}}
\newbox\bbox
\setbox\bbox=\hbox{{\bsy\char'015}}
\def\grin{\hbox{\box\bbox\kern-5.75mm\box\sbox}}

There are in fact plenty of smileys around, apparently because the IBM PC character set has two, but my closest was in the Metafoundry’s Hodge-Podge (mdhp) font, which would make TTN non-portable (sigh).

An award?

A plea to everyone reading this to help me with a new task. I want to collect samples of truly appalling desktop publishing for a booklet on typography. I mentioned this on the networks some while back, but have only had one or two contributions. You must all recognise the stuff when you see it: dozens of typefaces, columns skewed, wrong fonts, bad or no hyphenation … the things which make your stomach churn. If you come across this valuable rubbish, please send it to me, at the Computer Centre, University College, Cork, Ireland. Originals if possible, good photocopies otherwise, and if the document doesn’t identify the creator or people responsible, see if you can find out for me: I may have some tricky copyright negotiations, but I don’t want to change names to protect the guilty. Perhaps some rich person can institute an annual award for the Most Ghastly Piece of Typographic Rubbish. Thank you in advance.

“Hey — it works!”

Dollars and pounds, TEX and L^{A}TEX

Christina Thiele

In TTN 1,3, Alan Hoenig’s report on the Prague Meeting required the use of the pound sterling sign. I discovered something interesting when it comes to L^{A}TEX: the TEX strategy of typing {\it\$} to yield a pound sign (albeit an italic one), doesn’t work: you get what you coded for — $, an italic dollar sign! Hmmm … “When in doubt, read the documentation” — Lamport manual, p. 40, in Section 3.2.2. Symbols (and again on p. 157): \pounds will give you £.

But TEX or L^{A}TEX, the use of an italic pound sign along with a roman dollar sign bothers me.^{1} Quite a while ago, when I was printing up font charts, in order to see what was available, I ran across the cmu font: Computer Modern Unslanted. Check your collection of font files to see if you have cmu10; if you do, then you have an upright pound sign ($) at your disposal:

definition:  coding:

           In L^{A}TEX: \newfont{\unsl}{cmu10}  {\unsl\$}
           In TEX: \font\unsl=cmu10        {\unsl\$}

But what if you really do need the two signs in italics?! Well, that’s no problem for L^{A}TEX: {\it\$} gives you $, and \pounds gives you £; but if you’re using TEX, you need to do some more defining. There’s no real italic dollar sign, but the slant font cmsl10 has an adequate substitute: $. Comparing the slanted and italic dollar signs ($ vs. $), I find it hard to see a difference. … or is someone going to tell me that’s what Lamport did all along ;-)

I’ve checked with Johannes Braams about NFSS users: you don’t have to worry at all! The two signs will be set in either roman or italic automatically, depending on the context.

If you really want to know what’s under the hood of your TEX-mobile, find yourself a macro file for printing up font tables — if you don’t have e-mail access, write to me directly (address in the membership directory). If you do have net access, mail me at cthiele@ccs.carleton.ca.^{2} Print up everything you’ve got, and examine especially the lesser-used fonts such as cmu or dunhill or cmfib — it’s quite interesting what’s available. For more information on font loading and such, check the entries in the TEX and L^{A}TEX manuals, under font.

1I posted the query on mixing upright dollar signs with italic pound signs to the typo-l list (mentioned in Peter Flynn’s article in TTN 1,2) and it would seem that upright pound signs are more appropriate in regular roman text.

2I’ve only got a TEX file; anyone got a L^{A}TEX one they could offer?

(L^{A})TEX News

DC fonts — questions and answers (I)

Yannis Haralambous

                                    yannis@gat.citilille.fr

The following questions and answers are provided to start you thinking about DC fonts: what they are, why they exist, what benefits they provide, and so on. They are aimed at providing information from the user’s perspective, and so are intended to be pragmatic rather than theoretical. This is Part I; Part II will appear in the next issue of TTN.

Q: What are DC fonts? What makes them different from CM fonts? A: We should distinguish between two things: (a) fonts using the Cork Encoding scheme, and (b) DC fonts. Any font can be re-encoded according to the Cork Encoding;^{1} DC fonts are Computer Modern fonts, encoded this way — and since the Cork encoding involves characters not existing in the CM encoding, these have been designed from scratch by Norbert Schwarz. Here are some comments:

(a) The Cork encoding is the new standard encoding for fonts used by TEX; it bears a strong resemblance to ISO Latin 1; combined with TEX’s 3.0 multilingual features, it allows efficient typesetting in most European languages. An important difference between CM and Cork encodings is the fact that while the former involves 128 characters, the latter provides 255 characters (and one empty location for special purposes). This is the highest possible number of characters for a font used by TEX;^{2} these characters have been chosen so that the maximum number of languages is covered. Another difference is that the Cork font tables for roman, italic, typewriter, and small capital fonts are exactly the same.^{3} This makes Cork-encoded fonts more consistent than CM fonts.

(b) The shapes of unaccented characters in the DC fonts are identical to those of Computer Modern fonts; on the other hand, the heights of accents are quite different from those obtained by using the \accent primitive. An important new feature of DC fonts is that you can generate fractional point sizes. Norbert Schwarz has included an interpolation mechanism in the DC font package which allows to produce, for example, a 10.9545 point bold extended font out of the parameter values given by Knuth for the point sizes of the cmbx font.

1Defined at the 1990 TUG meeting, held in Cork. See Ferguson, “Report on Multilingual Activities”, TUGboat 11, no. 4, pp. 514–516.

2No doubt about it, a UNICODE version of some descendent of TEX will be accompanied by a new encoding standard, involving 2^{16} characters; but that program will not be called TEX, and the encoding will not be a “TEX standard”.

3The ligatures fi, ff, etc. are provided in typewriter form; however, they are not inserted automatically.

Q: What kinds of situations require DC fonts? What kinds of working environments would benefit right away from installing them? A: You need Cork-encoded fonts to typeset in any Latin-alphabet language, except English and Latin. With CM-encoded fonts you will have neither hyphenation nor correct upper- or lowercase forms. All working environments will benefit right away from installing them, except perhaps an hypothetical environment which

(a) uses TEX to typeset exclusively in English and/or Latin;

(b) has no connection with environments not satisfying condition (a);

(c) will continuously fulfill conditions (a) and (b) for a long period.

Q: If I’m using TEX in a completely English-language environment, do I need DC fonts? I don’t expect to receive foreign-language materials at all. A: In other terms, you satisfy conditions (a), (b) and (c) of the previous question. But nevertheless, you are now reading TTN, which means that you are concerned about new developments in TEX, and that you feel part of the big TEX community.

Allow me to give you another — less technical — reason for switching to Cork-encoded fonts: a consequence of Knuth’s announcement that there will be no further features added to TEX, is that he is not going to change the encoding of CM fonts, described in volumes A, C and E of Computers and Typesetting. While unfortunate, the fact is that this encoding does not allow typesetting in European languages. Somebody had to introduce a new encoding; the decision was taken during a meeting of the TEX Users Group (Cork 1990), after months of discussions and preparation. By adopting this new standard, we, the TEX community, allow TEX to be efficiently and consistently used by thousands of other people, and by this give it another chance to survive (remember, the big brothers are watching us …)

Q: Do you have to have TEX v.3.xx running? Our site only has 2.1. A: You need TEX v.3.xx for multilingual texts (because it allows you to use several sets of hyphenation patterns in the same text), and for certain scripts involving special ligatures (such as Greek, Hebrew or Arabic fonts). Corkencoded fonts can be used with previous versions of TEX. By the way, don’t you think it is time to upgrade?

Q: Where do I get them?

A: In theory they should be included in every TEX distribution, free or commercial. If you don’t have them, you can fetch them by anonymous ftp from

   the Stuttgart server: rusmv1.rus.uni-stuttgart.de (IP 129.69.1.12), file
   /soft/tex/fonts/metafont/dc.

Q: Do I access them in the same way as regular TEX and L^{A}TEX fonts? A: Yes, mainly because they are regular TEX and L^{A}TEX fonts. Actually, once installed in your system, you shouldn’t have to think about accessing them at all. Commands like \rm, \bf, or L^{A}TEX environments like small, verbatim will automatically call DC fonts instead of CM ones.

Lollipop now available _{Victor Eijkhout}

                                               eijkhout@cs.utk.edu

We all know that in TEX any layout is programmable. Unfortunately, for the people who actually know what to program, TEX macro writing is usually too hard. In fact, it is often hard for seasoned programmers. The Lollipop format tries to bridge this gap.

Lollipop is a macro package that gives the user powerful tools for programming macros for lists, headings, output routines and a number of other things. The Lollipop macros take the specifications for a macro, and then construct that macro: it is a package for writing macros.^{1}

About this release

Lollipop is not finished (the first public release was in October 1992, the current version 0.93 is dated November 1992), but it is already quite powerful.^{2}

Since Lollipop is still under development I will be very accessible for questions, bug reports and suggestions — e-mail is preferred. And I will be flattered if you actually use Lollipop, but beware that changes may occur in future versions.

Lollipop can be ftp’d from the directory /pub/eijkhout/tex on cs.utk.edu (log in as ‘anonymous’), where it is stored as lollipop.tar.Z. Uncompress and tar -fx it. It is also available by e-mail from listserv@hearn.bitnet; send

the message: GET LOLLIPOP PACKAGE.

Like a lot of TEXware, Lollipop is free. Make and give away all the copies that you want. Just don’t ask money for it, other than reasonable expenses such as for duplicating discs or photocopying the manual. Share and enjoy!

1[Lollipop was also the subject of a paper presented at the TUG’92 meeting in Portland, Oregon; see TUGboat 13, no. 3, pp. 341–346 –Ed.]

2It was used to typeset my book TEX by Topic, for instance. I suggest that anyone who is interested should format the manual and have a look at the examples in it. They document the current power of Lollipop.

NTG’s MAPS bibliography database now available

Gerard van Nes

                                                     vannes@ecn.nl

MAPS (Minutes and APpendiceS) is the informal newsletter of the Dutch TEX User Group (NTG). Starting with 1992, the corresponding maps.bib file is available by anonymous ftp (and e-mail) from archive.cs.ruu.nl in directory

pub/tex/ntg, or by e-mail from the NTG fileserver tex-nl@nic.surfnet.nl.

As most readers of TTN know, the TUGboat bibliography file tugboat.bib, covering the years 1980–1992, is already available by anonymous ftp to

ftp.math.utah.edu in directory pub/tex/pub/tugboat. It can also be found

with a large collection of other bibliographies and bibliography tools in

pub/tex/bib.

Other Local User Groups are urgently requested to make their own .bib files also available to the whole TEX community. Including abstract fields would be especially welcome in order to inform interested users about the contents of the local articles. Ftp and/or e-mail addresses for the local .bib files should be forwarded to the editor of TTN (cthiele@ccs.carleton.ca). Examples of the bibliography files can be found in tugboat.bib or maps.bib.

TUGlib archive — update

Nelson H.F. Beebe University of Utah

                                               beebe@math.utah.edu

The TUGlib server on math.utah.edu provides electronic mail access to the TUGlib archives. Besides giving access to a large collection of TEX files and support software, this server makes the entire anonymous ftp collection on ftp.math.utah.edu available as well. This includes software for UNIX, VAX VMS, and the IBM PC.

If you have Internet anonymous ftp capability, you should always use that in preference to electronic mail, because it is faster and much more convenient. Sun workstation users may enjoy ftptool, available via anonymous ftp to sunsite.unc.edu from the directory pub/X11/Openlook. At the time of writing, the current version of ftptool is in the file Ftptool4.3.tar.Z. Ftptool provides a graphical user interface to the ftp program, with point-and-click operations for login, directory listings, moving around in the directory tree, viewing of documentation files, and file transfers.

   The anonymous ftp server on ftp.math.utah.edu is a new program that

offers the ability to fetch an entire subtree in UNIX tar, or in compressed tar.Z, Info-ZIP zip, and Dhesi zoo formats. The compression reduces the network transfer overhead, while the archive formats preserve file protections and time stamps, and support a file directory hierarchy. The top-level README file provides further details, and the IAFA* files follow Internet anonymous ftp archive guidelines for providing information about the archives.

If you only have electronic mail access, then to get started with TUGlib, send a message with the text help to tuglib@math.utah.edu. This will return enough information for you to understand how the server works. A typical request for the index to a directory is send index from tex/bib, and a request for a specific file, send tugboat.bib from tex/bib. Several requests can be sent in a single message.

TUGlib also supports queries of the TUG membership list; for details, consult the help response message.

New list for METAFONT users and afficionados

Yannis Haralambous

                                           yannis@gat.citilille.fr

The programming language is a faithful companion to TEX. To use TEX, you must have some notions of it; this is not the case for . Many people use it inside batch mode processes to create missing fonts, without ever having looked inside a .mf file. Other people — adventurers — have ignored Knuth’s warning on p. viii of the book, and have inevitably become hooked by . Thanks to the French-speaking TEX Users Group GUTenberg a new e-mail discussion list has been created,

1. as a means of communication between hooked ers;

2. as a way to bring the ‘rest of us’ closer to them;

3. as a means to get quick and efficient answers to questions such as

• why do I always get a .2602gf file?

• what is a “strange path” and what can I do to avoid it?

• is there a way to go from to PostScript and vice-versa?

• where can I find a Stempel Garamond font written in ?

• what is meta-ness?

4. and finally, as a first step to encourage people to undertake ing, and start a new post-Computer Modern era of !

To subscribe to this list, send the following two lines to listserv@ens.fr on the Internet:

SUBSCRIBE METAFONT <Your name and affiliation>

   SET METAFONT MAIL ACK

The address of the list is metafont@ens.fr (at the notorious École Normale Supérieure de Paris). Owner of the list is Jacques Beigbeder (beig@ens.fr), coordinator is Yannis Haralambous (yannis@gat.citilille.fr). Language of the list is English; intelligent mottos are encouraged. FAQ files will be written from time to time; their location and contents will be announced through the list. Come and join us!

News from Around …

Electronic documents and the law _{Jacques André}

                                        Jacques.Andre@IRISA.FR

AFCET’s “Technical Committee on Electronic Documents” is organising a 1-day meeting (for some time in the Spring) on the issue of electronic documents and the law in France. The Committee is looking for questions and topics which computer scientists, specialists in the management of electronic documents, editors, word processing personnel, and others would likely ask. Some sample situations:

• Can I copy a friend’s fonts and use them: for personal stuff (e.g., correspondence)? to produce a publication which is free? to produce a book for sale?

• I’ve written a book and provided the publisher with a disc containing both the source file and the style file (e.g., DSSSL, L^{A}TEX or Framemaker, a style sheet for Word, etc.).

– Do I retain copyright over the style, and if so, on what grounds?

– Can the publisher re-use my style file for another book?

– Let’s say that I have some original things in the style file. How can I prevent the editor from reading it (and eventually copying it)?

• I’ve bought a hypertext disc on twentieth-century painting. Can I delete the files on it and replace them with my own files on, say, grasshoppers and distribute it to my students taking African ecology?

• Can I sell a syntax checker which:

– uses a database I’ve built by using a CD-ROM version of the Oxford English Dictionary (OED)?

– directly access the OED on CD-ROM, which I require potential clients to purchase separately?

Send your questions and problem situations — even if you think it’s too old-hat or silly — to:

Jacques André  e-mail: jandre@irisa.fr
IRISA/INRIA-Rennes  Tel: (33) 99 84 73 50
Campus de Beaulieu  FAX: (33) 99 38 38 32

F-35042 Rennes cedex, France

Reports on Meetings

UKTEXUG: 14 October AGM _{R.A. Bailey}

UK TEX Users’ Group

At the recent Annual General Meeting of the UK TEX Users’ Group, held on October 14th, elections were held and the UKTUG committee now consists of P. Abbott (Chair), R.A. Bailey (Committee Secretary), S. Brooks (Baskerville editor), D. Eckersley, R. Fairbairns, I.W. Hall (Treasurer), D. Murphy (Membership Secretary), D. Penfold (Membership Secretary), J. Petts, S.P.Q. Rahtz, P. Taylor (Acting Chair), and G. Toal.

After the AGM there was an ordinary meeting entitled “Something completely different and yet integrated”. R.A. Bailey, C.A. Rowley and D.V.J. Murphy spoke about using TEX, L^{A}TEX, AMS-TEX, AMS-L^{A}TEX and LAMS-TEX to wordprocess and typeset mathematical documents. After tea, D. Osborne showed participants how to access the Aston Archive using Internet tools.

The UK TEX Users’ Group will be establishing a fund to be called the Cathy Booth Memorial Fund. The Fund will be used to support education and research in electronic publishing in general and in the use and development of TEX and its relatives in particular, and for other charitable purposes concerned with education. UKTUG will be requesting donations to the CBMF some time within the next six months.

Upcoming meetings are tentatively planned for January or February, March, and May (details to be announced later).

NTG in 1993: 11th Meeting _{Gerard van Nes}

Secretary, NTG

The 11th Dutch TEX Users Group Meeting will be held on June 10, 1993 at the Royal Dutch Meteorological Institute in the heart of Holland: De Bilt.

The topic of this one-day meeting will be “From Font To Book”. Rein Smedinga (Computer Science Groningen) will explain how to create a book using L^{A}TEX. Physicist Lambrecht Kok (director of Coulomb Press Leiden) will speak about his publishing experience.

In addition, famous font designers will discuss all the in’s and out’s of fonts. NTG president Kees van der Laan will outline the history of the NTG’s five years of existence, and will look into the future.

For detailed information, contact Theo Jurriens, taj@astro.rug.nl.

TUG Board Activities

New TUG Executive Director named _{Malcolm Clark}

TUG President

Patricia Monohon has been appointed as TUG’s new Executive Director. She takes over from Ron Whitney, who has served TUG since 1989, first as production assistant for TUGboat, and then as TUG’s Business Manager. His plans for the future include consulting, freelance programming, and raising a child.

Pat is currently a TEX consultant and supports the Center for Remote Sensing and Environmental Optics at the University of California, Santa Barbara. She was instrumental in the formation of the Santa Barbara Area TEX Users Group. She currently teaches classes in both TEX and L^{A}TEX on the UCSB campus. Before this she was employed at the European Organization for Nuclear Research (CERN) in Geneva, where she installed TEX on a wide range of platforms, and provided a variety of other support roles, from the creation of macros through to instruction of personnel.

She also spent some time at the International Institute for Applied Systems in Laxenburg, Austria, where she edited conference proceedings, research papers and books, activities which she has since pursued in a freelance capacity. Prior to that, Pat was administrative services manager in the Math Dept. at the University of Washington.

She therefore comes to TUG with a proven record in administration, TEX support, the organisation of volunteers, and the editing and typesetting of books and journals. In addition, Pat has a working knowledge of French and German, a by-product of her years spent in Europe.

Pat is well known to many in the TEX world. She has been using TEX since 1982, and became a member of TUG in 1984. She has attended many of the annual TUG conferences held in Europe. I am very pleased to welcome her aboard.

A new TUG President, too _{Malcolm Clark}

TUG President

Christina Thiele has been elected President of the TEX Users Group. She will take over on January 1st, 1993, for a two-and-a-half year term of office.

This was a somewhat unusual election. You will probably have realised that you did not receive a list of candidates and a ballot slip. The intended procedure was that candidates would be nominated from the membership, and a ‘normal’ democratic postal voting system would swing into action. Sadly, by the due date, no candidates had appeared. This was quite a worry, especially to me, since I might conceivably have to continue longer as President, something I had not bargained for. We might argue that the Vice President would take over on January 1st, but Ken Dreyhaupt was unable to take on this position. We were therefore left in a rather awkward position. Exercising our executive powers, the Executive Committee devised the following scheme: we would, in the first instance, invite members of the Board to put themselves forward for election to the office; in so doing we stressed the importance of the post, the fact that they were already elected members, and that they had been elected to serve the membership — in other words, we appealed to their better nature and honour; failing this we would have coerced, cajoled or conscripted more widely.

Fortunately there were two honourable Board members who felt the call to serve. I don’t want to make light of this: not every Board member is in a position to take on the role of President, but their contribution on the Board may be very valuable and significant. Both Christina Thiele and Mimi Jett stepped forward. We then circulated ‘position statements’ or ‘manifestos’ from the two candidates to all members of the Board, and voted, by e-mail, with the result that Christina was elected President.

Why didn’t we circulate this information to the membership, and proceed with a more normal election? We were running out of time. At the time it was evident that no candidates were forthcoming, we had no contingency plan — something the election procedures committee will be examining in the future. Why there should have been no candidates is another question — in all modesty I assume it is because I am such a hard act to follow. The time taken to discuss and circulate a plausible substitute procedure, then find a volunteer, and provide a decent period for voting, meant that we were close to the end of November before the votes could be collected and counted. Had we decided to adopt a postal vote of all members, we might easily have been well into January before the results could be announced.

I don’t pretend that this is the best possible procedure which can have been erected. I think however it was a reasonably fair procedure. I do regret that we were left with insufficient time to advise the whole membership of the activities. But there are times when leadership must be exercised. I am particularly encouraged that I, for one, would have been happy to see either candidate as President. They both have proven, excellent records in the group, and outside.

We have a new, capable President, and we have a new capable Executive Director. They are both women: things are definitely improving!

TEX Users Group 1993 Course Schedule

Beginning/Intermediate TEX* Intensive Course in L^{A}TEX*
Boston  April 19–23  San Francisco March 1–5
San Diego  June 14–18  Boston  April 26–30
Boston  August 9–13  Ottawa  August 23–27
Chicago  Oct. 18–22  Boston  Oct. 25–29
Modifying L^{A}TEX Style Files* Adv’d TEX and Macro Writing*
San Diego  June 7–11  San Diego  June 21–25
Boston  Nov. 1–5
TEX for Publishers  Practical SGML and TEX*
San Francisco  March 8  San Francisco March 9–10
Boston  March 22  Boston  Nov. 8–9
New York City June 9
Washington, DC Nov. 12
Book and Document Design  SGML and TEX for
with TEX  Publishers*
Boston Sept. 23–24  New York Nov. 10

*Lab classes — computers will be provided for all students.

• TUG courses are small, with 8–15 students in most classes, and are held at major hotels.

• The dates and locations above are tentative — watch for the final schedule in the mail this fall.

• For more information, contact the TEX Users Group at 805-899-4673.

On-Site Courses in TEX and L^{A}TEX from TEX Users Group

• Courses in TEX, L^{A}TEX, SGML and TEX, PostScript, or TEX for Publishers tailored to the needs of your group

• Courses at every level from beginning to advanced

• Five full days of instruction at your site

• One-week course fee includes all instructor fees and expenses plus textbooks and other materials for up to 15 students

• If a properly equipped training facility is not available, TUG will arrange computer rentals and supply TEX or L^{A}TEX software

Upcoming Events

{24–27 Feb. \atop 1993}

CONCEPPTS’93:

The Prepublishing Conference Orange Co. Convention Centre, Orlando, Florida.

Phone: 703-264-7200 FAX: 703-620-9187.

1 March Knuth Scholarship: Deadline for 1993 submissions. (see Cover 2 for address)

Knuth Scholar. Committee c/o TUG Office

                           tug@math.ams.org

9–12 March DANTE’93 and General Meeting, Chemnitz (near Dresden).

Dr. Wolfgang Riedel

                           wolfgang.riedel@hrz.tu-
                           chemnitz.de

30–31 March

UKTUG:

“TEX for non-American languages”. Conference and tutorials, University of Glasgow Phil Taylor

                           chaa006@vax.rhbnc.ac.uk

10 June NTG: 11th Annual Meeting, Royal Dutch Meteorological Institute, De Bilt.

Theo Jurriens

                           taj@astro.rug.nl

16–19 June

ACH-ALLC93:

Joint int’l. conference: Assoc. for Computers and the Humanities and the Assoc. for Literary and Linguistic Computing Georgetown University, Washington, D.C.

Dr. Michael Neuman

                           Neuman@Guvax.bitnet
                           Neuman@Guvax.Georgetown.
                           edu

26–29 July

TUG’93:

“World Wide Window on TEX” Aston University, Birmingham, UK.

Chris Rowley, ca rowley

                           @vax.acs.open.ac.uk

Malcolm Clark

                           malcolmc@wmin.ac.uk

Note: Also consult the “Calendar” in the previous issue of TUGboat for more dates.

Index

TTN 0,0 to 1,4 (inclusive)

This Index was prepared by Jackie Damrau; it includes material which has appeared in 5 issues of TTN, from the prototype of 1991 to this, the final issue for 1992. Notation: All references show the volume,number:page; outside back cover is noted as C4.

– C –

1993 Course Schedule 1,3:26; 1,4:24

– E –

Editorials 0,0:1; 1,1:1; 1,2:1; 1,3:1; 1,4:1; Elections

Board 0,0:25; 1,1:2 President 1,2:28; 1,3:25; 1,4:22

Electronic Retrieval Sites New list for METAFONT users 1,4:19 New LISTSERV for Greek TEX Users 0,0:6 DC fonts 1,4:17

– F –

Feature Articles

jeep.sty: A General Purpose Style File 0,0:3

Scientific Publishing in Mexico 1,1:7 The Technical Council 1,3:5 TEX Conventions Concerning Languages

1,4:3 Update on TEX in Japan 1,2:5

– H –

“Hey — it works!”

Dealing with very long paragraphs 1,1:15 Dollars and pounds, TEX and L^{A}TEX 1,4:14 Some applications for \llap 0,0:30 The Cauchy integral 1,1:15; 1,2:15 Turning double brackets into delims. 1,2:14 Marking lines in the margin 1,3:12; 14:10

Editor’s note 1,3:14 When there’s only one of you … 1,3:14 Some TTN L^{A}TEX tricks 1,3:15

– I –

In Memoriam Cathy Booth 1,1:28 Samuel B. Whidden 1,1:28

Index 1,4:26

– K –

Knuth Scholarship

1992 1,1:3; 1,3:3; 1,4:28

– L –

(L^{A})TEX News

DC fonts — questions and answers (I) 1,4:15

IL^{A}TEX1,1:16 L^{A}TEX Tables: workshop handouts 1,2:16 Lollipop now available 1,4:17 New list for METAFONT afficionados 1,4:19 NTG’s MAPS bibliography database 1,4:18 Recent Updates to L^{A}TEX 2.09 1,2:17 TEX Macro Index 1,2:16 TUGlib archive — update 1,4:18 Volunteer Work for the L^{A}TEX3 Project

1,3:18 L^{A}TEX3 Project

Bowling Tournament Boosts Fund 1,3:21 L^{A}TEX3, TUG, and You 1,2:2 L^{A}TEX3 Update 1,3:4 The L^{A}TEX3 Project Fund 1,1:5 Volunteer Work for the L^{A}TEX3 Project

1,3:18 What is the L^{A}TEX3 Project? 1,1:4

– M –

Meeting Announcements EuroTEX’92 1,2:22 TeCH’92 1,1:21 TUG’91: TEX Inroads into Publishing 0,0:28 TUG’92: TEX in Context 1,1:22, C4 TUG’93: World Wide Window on TEX 1,4:C4

– N –

New Publications 1,1:14; 1,2:13; 1,3:9; 1,4:11 News from Around …

CyrTUG 1,1:17 Electronic documents and the law 1,4:20 Estonian User Group 1,2:19 Gleanings From Here and There 1,3:19 GUST 1,2:19 GUTenberg 1,2:19 New LISTSERV for Greek TEX Users 0,0:6 SibTUG 1,1:18 TEX in the Soviet Union 0,0:5

– P –

Published TEX Credits

The Truth about TEX 0,0:6 Publications

EuroTEX’92 Proceedings (Prague) 1,3:28 New Publications 1,1:14; 1,2:13; 1,3:9;

1,4:11 NTG’s MAPS bibliography available 1,4:18 TEXniques (latest issues) 1,3:C4 TUGlib archive — Update 1,4:18

– R –

Reports on Meetings CsTUG 1,1:20 EuroTEX’91 1,1:20 EuroTEX’92 1,2:22; 1,3:22

EuroTEX’92 Proceedings (Prague) 1,3:28 Grupo de Usuarios de TEX 1,2:21 NTG Spring Meeting 1,2:22 NTG in 1993 1,4:21 TUG’91 1,1:19 TUG’92 1,3:19

Bowling Boosts L^{A}TEX3 Fund 1,3:21 UKTEXUG 1,2:20; 1,4:21

– S –

Style Files

                                           jeep.sty 0,0:3

– T –

The Technical Council 1,3:5 see also: Working Groups

TEXhax Tidbits 0,0:27

TEXniques (latest issues) 1,3:C4

TUG Board Activities Communications 0,0:20 Elections

Board 0,0:25; 1,1:2

President 1,2:28; 1,3:25; 1,4:22 Message from the Board 0,0:7 New TUG Executive Director named 1,4:22 A New TUG President, too 1,4:22 Summaries of Meetings

1990 0,0:11

1991 0,0:15; 1,1:23, 25 1992 1,2:24; 1,3:23

Typographer’s Inn 1,1:12; 1,2:11; 1,3:10; 1,4:12

– U –

Upcoming Events 0,0:32; 1,1:27; 1,2:27; 1,3:27; 1,4:25

– W –

Working Groups (WG-92-00) 1,3:6 (WG-92-01) 1,3:7 (WG-92-02) 1,3:7 (WG-92-03) 1,3:7; 1,4:3 (WG-92-04) 1,3:8 (WG-92-05) 1,3:8

Attention: Submissions for Knuth Scholarship Competition

Nico Poppelier

Board Liaison, Knuth Scholarship Committee

The TUG Office has moved to a new location, namely to Santa Barbara. Applicants should therefore request information and send submissions to the new address — see inside front cover. For more information on the scholarship, please read the announcement in TUGboat 13, no. 3, pp. 395–396.

I should like to stress that the competition is open only to members of the TEX Users Group with qualifications as explained in that announcement.

New TUG Address

TEX Users Group P.O. Box 869 Santa Barbara, CA 93102 USA Balboa Building, Room 307 735 State Street Santa Barbara, CA 93101 USA Phone: 805-899-4673

                      E-mail: tug@math.ams.org

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 TEX Conventions Concerning Languages

Yannis Haralambous . . . . . . . . . . . . . . . . . . . . . . . 3 New Publications

Peter Schmitt . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Typographer’s Inn

Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 “Hey — it works!” . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Dollars and Pounds, TEX and L^{A}TEX Christina Thiele . . . . . 14 (L^{A})TEX News . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

DC Fonts: Questions and Answers (I) Yannis Haralambous . . 15 Lollipop now available Victor Eijkhout . . . . . . . . . . . . . . 17 NTG’s MAPS database now available Gerard van Nes . . . . . 18 TUGlib archive — update Nelson H.F. Beebe . . . . . . . . . . 18 New list for METAFONT afficionados Yannis Haralambous . . 19

News from Around . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Electronic documents and the law Jacques André . . . . . . . . 20 Reports on Meetings . . . . . . . . . . . . . . . . . . . . . . . . . . . .

UKTEXUG: 14 October AGM R.A. Bailey . . . . . . . . . . . 21 NTG in 1993: 11th Meeting Gerard van Nes . . . . . . . . . . 21

TUG Board Activities . . . . . . . . . . . . . . . . . . . . . . . . . . . New TUG Executive Director named . . . . . . . . . . . . . . . . 22 A new TUG President, too . . . . . . . . . . . . . . . . . . . . . . 22

TEX Users Group 1993 Course Schedule . . . . . . . . . . . . . . . . . 24 Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Index (TTN 0,0 to 1,4, inclusive) . . . . . . . . . . . . . . . . . . . . . 26 Attention: Submissions for Knuth Scholarship Competition . . . . . . 28

Volume 1, No. 4, 1992

TUG93 call for papers

World Wide Window on TEX

14th Annual TEX Users Group Meeting July 26th – 29th, 1993

Aston University in Birmingham, UK, will be the venue for the 1993 TUG conference. Aston is the home of the Aston Archive, one of the largest collections of electronic TEX paraphernalia. This is the first time the annual meeting will have been held outside North America. The location of the conference at one centre of the electronic web and its movement from North America encourages particular focus on the world-wide aspects of TEX (L^{A}TEX, , …). The marked rise in maturity of windowing systems (Macintosh, Atari, Amiga, Windows3, X windows) also allows us to exploit more straightforward and direct ways of employing the TEX tools. It is hoped that there will be a contribution to the conference from the Didot project, further extending the range of topics to include digital typography and font creation. The conference will feature the regular paper presentations, but workshops, poster displays, courses, panels and birds-of-a-feather sessions will be integral components.

Contributions are being actively sought in the following subject areas: ⋄ archives ⋄ electronic networks ⋄ formatting structured documents ⋄ L^{A}TEX3 ⋄ graphical user interfaces to TEXware ⋄ non-english issues ⋄ non-Latin scripts ⋄ digital typography ⋄ editing structured documents ⋄ styles ⋄ other typesetting systems ⋄ document views ⋄

Abstracts should be sent to the program coordinators by February 22nd, 1993.

Conference committee Maureen Campbell, Peter Abbott, Chris Rowley, Philip Taylor, Carol Hewlett, Sebastian Rahtz, Malcolm Clark

Program coordinators

Chris Rowley  Malcolm Clark
Parsifal College  IRS
Open University  University of Westminster
Finchley Road  115 New Cavendish Street
London NW3 7BG  London W1M 8JS
phone: 071 794 0575  071 911 5000 ex 3622

 e-mail: ca_rowley@vax.acs.open.ac.uk  malcolmc@wmin.ac.uk

fax:  071 433 6196  071 911 5093